The Thai Baht fell by 17 against the US Dollar. by how much has the US Dollar appreciated against the Baht?

please i need an answer to this with steps
my teacher put this solution but i still can't understand how we got to this answer so can anyone provide me with a different solution please?

e-eo co o,33 ot|+ Q83eo eo 193eo 183eo ( 83) eo во = 248 - 20,95.
Answer #1

The solution you have provided by your professor is the standard form for calculating what have been asked in the question

i will give you another method of doing it hope you will understand

so according to the question thai fell 17 against the dollar

Now suppose the value of dollar is 100 so the value of thai = 100-17= 83

this means that 83 is 17% less than 100

what we have to calculate is that how much more 100 is that of 83

in simple words we can say that what percentage of 83 is added in 83 so that we can get 100

putting this in equation we get

83 + 83*y/100 = 100 (y= rate or percentage)

83/1 + 83 y/100 =100 ( taking L.C.M)

8300 + 83y/ 100 =100

8300 + 83y = 100* 100

8300 + 83y = 10000

83 y = 10000 - 8300

83y = 1700

y= 1700/83

y= 20.481

so dollar has appreciated 20.481 against baht